OpenStudy (anonymous):

Simplify the radical. sqrt 200 would it be 2 sqrt 10 ?

OpenStudy (anonymous):

10 sqrt 2

OpenStudy (anonymous):

sqrt(200) sqrt(2*100) sqrt(2*10^2) 10sqrt(2)

OpenStudy (callisto):

Similar to your previous one \[\sqrt{200}=\sqrt{2 \times 10 \times 10}=\sqrt{2 \times 10^2} =\sqrt{2} \times \sqrt{10^2} = \sqrt{2} \times 10 =10\sqrt{2}\]

OpenStudy (muhammad_nauman_umair):

10sqrt(2)